The thigh has a mass of 4kg and a location of (20,18). The shank has a mass of 2.5kg and a location of (24,15). The foot has a mass of 0.8 kg and a location of (30,10).

What is the Center of Mass for the leg?

Question 9 options:

A)

(23.5, 15.4)

B)

(25, 14)

C)

(24,15)

D)

(22.5, 16.1)
Answer #1

***********************************************************************************************
Check the answer and let me know whether it's correct.....
Answer #2

SOLUTION :


Let centre of mass for the leg be at point (x, y)


Total mass  = 4 + 2.5 + 0.8 = 7.3 kg


So,

 

7.3 * x = 4*20 + 2.5*24 + 0.8* 30 = 164

=> x = 164/7.3 = 22.47 


7.3 * y = 4*18 + 2.5*15 + 0.8*10 = 117.5

=> y = 117.5/7.3 = 16.10 


Hence,


Center of mass for the leg is at point (22.47, 16.10) : Option D (ANSWER).

answered by: Tulsiram Garg